Linear systems:
2x-3y=0
4x-6y=0

Answer by venugopalramana(3286)

SEE THE FOLLOWING EXAMPLE AND DO THE SAME WAY>>>
2x-3y=5
4x-6y=10
the second eqn. is obtainedby multiplying with 2 the first eqn.
so they are dependent eqns..so in effect we have only one eqn.with 2 variables which give raise to infinte solutions
let us take 2x-3y=5
put x=0,1,2,3,etc..and get corresponding values of y
x..0......1......2...3...4...5..
y..-5/3...-1...-1/3.....etc..infinite solutions
.